OpenStudy (anonymous):

Helpp? If a tennis ball has a diameter of approximately 6.5 centimeters, what is its surface area?

OpenStudy (anonymous):

What's the equation for finding the surface area of the sphere?

OpenStudy (anonymous):

radius is half the diameter --> 1/2*6.5 = 3.25 SA = 4*pi*(3.25)^2 pi is 3.14 SA = 4(3.14)(3.25)^2 = 132.7

OpenStudy (anonymous):

SA = 4pi*r^2

OpenStudy (anonymous):

Make sense?

OpenStudy (anonymous):

@IndianChick500 ?